# JailBound: Jailbreaking Internal Safety Boundaries of Vision-Language Models

### v2

@@ Abstract
Vision-Language Models (VLMs) exhibit impressive performance, yet the integration of powerful vision encoders has significantly broadened their attack surface, rendering them increasingly susceptible to jailbreak attacks. However, lacking well-defined attack objectives, existing jailbreak methods often struggle with gradient-based strategies prone to local optima and lacking precise directional guidance, and typically decouple visual and textual modalities, thereby limiting their effectiveness by neglecting crucial cross-modal interactions. Inspired by the Eliciting Latent Knowledge (ELK) framework, we posit that VLMs encode safety-relevant information within their internal fusion-layer representations, revealing an implicit safety decision boundary in the latent space. This motivates exploiting boundary to steer model behavior. Accordingly, we propose JailBound, a novel latent space jailbreak framework comprising two stages: (1) Safety Boundary Probing, which addresses the guidance issue by approximating decision boundary within fusion layer’s latent space, thereby identifying optimal perturbation directions towards the target region; and (2) Safety Boundary Crossing, which overcomes the limitations of decoupled approaches by jointly optimizing adversarial perturbations across both image and text inputs. This latter stage employs an innovative mechanism to steer the model’s internal state towards policy-violating outputs while maintaining cross-modal semantic consistency.
Extensive experiments on six diverse VLMs demonstrate JailBound’s efficacy, achieves 94.32% white-box and 67.28% black-box attack success averagely, which are 6.17% and 21.13% higher than SOTA methods, respectively. Our findings expose a overlooked safety risk in VLMs and highlight the urgent need for more robust defenses. Warning: This paper contains potentially sensitive, harmful and offensive content.
@@ 1 Introduction
Vision Language Models (VLMs), exemplified by GPT-4o [12], LLaVA [16], and Qwen-VL [1], have emerged as powerful agents capable of sophisticated multimodal reasoning. VLMs achieve their capabilities by integrating pre-trained vision encoders with Large Language Model (LLM) backbones [15], inheriting the strengths of both visual perception and natural language understanding.
Despite extensive safety alignment efforts in LLMs [13, 28, 10, 26] to mitigate harmful outputs, the application of similar safeguards to VLMs [19, 9] remains fraught with critical challenges. The integration of visual modalities significantly broadens the attack surface, making VLMs increasingly susceptible to jailbreak attacks that threaten their safe deployment [29, 11, 14, 30, 31]. Such attacks leverage adversarial perturbations, subtly introduced into either images [22, 2, 8, 35] or texts [32, 37, 17, 35, 25], to elicit harmful responses. Existing jailbreak strategies, explored across diverse white-box [31, 20, 2, 32, 8] and black-box [33, 5, 6, 7, 34] scenarios, highlight this vulnerability. However, these methods prove inadequate due to two primary deficiencies: 1) Depending on gradient-based optimization that are susceptible to be trapped in local optima, lacking precise directional guidance and resulting in suboptimal or overly conspicuous perturbations. 2) Processing visual and textual inputs separately, neglecting crucial cross-modal interactions.
@@
Intriguingly, recent works about the Eliciting Latent Knowledge (ELK) [3, 23] on LLMs, reveal a critical phenomenon: models often encode an implicit understanding of safe versus unsafe behaviors within their internal hidden states. This suggests the existence of an internal latent safety decision boundary. Inspired by these findings, we investigate whether this concept extends to the multimodal realm of VLMs. By identifying and manipulating the latent representations associated with an internal safety boundary, we can effectively trick the model into perceiving a harmful input as acceptable for output. Building on this insight, we propose a novel attack framework, JailBound, that targets VLMs’ internal decision boundaries. JailBound first probes to approximate this latent safety boundary, and then strategically manipulates cross-modal latent representations towards the target decision, thereby eliciting policy-violating responses, as shown in Figure 1.
@@ 2 Related Work
Eliciting Latent Knowledge.  Recent research increasingly focuses on extracting and understanding the internal knowledge representations within large language models (LLMs), often called Eliciting Latent Knowledge (ELK). This inquiry is motivated by the observation that LLMs may hold knowledge suppressed by safety alignment or training biases. A foundational work is raised by Christiano et al. [27], which formally posed the ELK problem. Building upon this, Burns et al. [4] introduce Contrast-Consistent Search (CCS), an unsupervised method for identifying truthful directions in activation space. CCS leverages the logical consistency properties of truth to identify these directions without using any labeled data or model outputs. Mallen et al. [24] extend the ELK framework with “quirky” LLMs and datasets to evaluate the robustness of ELK methods. Our work builds upon these foundations by leveraging the insights from ELK research to address the specific challenge of jailbreaking vision-language models (VLMs). We build on ELK insights to tackle jailbreaking in VLMs, hypothesizing that policy-violating knowledge may persist internally. By adapting and extending ELK methods, we aim to exploit this latent knowledge to circumvent safety mechanisms and elicit harmful or policy-violating responses.
Jailbreak Attacks on VLMs.  The integration of vision and language modalities in VLMs introduces new attack surfaces. Recent research [29, 6, 11, 14, 30, 31, 25, 21] have focused on jailbreak attacks, which aim to circumvent safety mechanisms and elicit harmful or policy-violating outputs from these models. For instance, FigStep [11] uses typographic images to evade text-based filters, while VAJM [29] shows adversarial images can bypass the safety mechanisms of VLMs, forcing universal harmful outputs. SCAV [32] manipulates latent embeddings within LLMs to elicit unsafe outputs, though conceptually similar in targeting internal representations, it is restricted to uni-modal settings and lacks the ability to manipulate visual content or transfer to black-box models. Other works [2, 8, 22, 35] explore adversarial perturbations, subtly modifying input images or text to mislead the VLM. These perturbations can be optimized using gradient-based methods in white-box settings [31, 20, 2, 32, 8] or through black-box approaches [33, 5, 6, 7, 11, 34] that rely on query feedback. Our work builds upon these advancements but is uniquely motivated by the principles of ELK, we aim to exploit this latent knowledge to craft jailbreak attacks that are both effective and semantically coherent. This approach allows us to probe the decision boundaries of VLMs, as informed by the latent representations, and generate perturbations that are more likely to succeed.
@@ 3.1 Problem Formulation
For the textual modality, An input text prompt is first processed by a tokenizer, which converts the raw character sequence into a sequence of token IDs, $X_{t}^{\text{raw}}=[t_{1},\ldots,t_{L}]$, where each $t_{i}\in\{1,\ldots,W\}$ and $W$ is the vocabulary size. These token IDs are subsequently mapped to dense vector representations by then embedding layer $f_{t}:\{1,\ldots,W\}^{L}\rightarrow\mathbb{R}^{L\times D_{e}}$, which produces the textual feature $x_{t}=f_{t}(X_{t}^{\text{raw}})\in\mathbb{R}^{L\times D_{e}}$. For textual attacks, we append a crafted suffix $X_{t}^{\text{suffix}}$ to form $X_{t}^{\text{adv}}=[X_{t}^{\text{raw}},X_{t}^{\text{suffix}}]$, and obtain perturbed features $\tilde{x}_{t}=f_{t}(X_{t}^{\text{adv}})$.
@@ 3.4 Joint Optimization Strategy
1:Input: Dataset MM-Safetybench $\mathcal{D}=\{(X_{v}^{(i)},X_{t}^{(i)},y^{(i)})\}_{i=1}^{N}$; VLM $F_{\theta}=(f_{v},f_{t},\phi)$; Target threshold $P_{0}$
@@
12:      Learn $(w^{(l)},b^{(l)})$ by minimizing the logistic regression loss in Eq. (3)
@@
1:$X_{v}^{\text{raw}}$, $X_{t}^{\text{raw}}$; VLM $F_{\theta}$, fusion layers $\{\phi^{(l)}\}_{l\in\mathcal{L}}$; step sizes $\eta_{v},\eta_{t}$; $(v^{(l)},\varepsilon^{(l)})$
@@ 4.1 Experimental Setup
To validate the efficacy of our JailBound framework, we conducted extensive experiments across six Vision-Language Models (VLMs), encompassing both white-box and black-box scenarios. We leveraged the MM-SafetyBench dataset [18], a meticulously curated multimodal safety evaluation benchmark. This dataset holistically covers 13 prohibited content categories, derived from prevailing AI safety policies and emergent multimodal threat vectors. Encompassing 1,719 adversarial examples across diverse risk scenarios, each sample pairs unsafe visual content with corresponding malicious prompts. This comprehensive curation ensures a robust evaluation of our method’s resilience against both conventional and emerging threats.
@@ 4.2 Comparison With Unimodal Attacks
We first evaluate the attack efficacy of JailBound in a white-box setting across Llama-3.2-11B, Qwen2.5-VL-7B [1], and MiniGPT-4 [36] in Table 2. As shown in Table 3. We use Attack Success Rate (ASR) as our evaluation metric, which refers to the proportion of successful adversarial attacks among all attempted attacks. For a more detailed discussion, please refer to Appendix. Our method demonstrates superior ASR across all models, achieving average ASRs of 94.38%, 91.40%, and 97.19%, respectively.
@@ 4.3 Comparison With Multimodal Attacks
To validate the effectiveness of our Decision Boundary-Aware Optimization, we compare JailBound against several existing attack methods, including UMK[31], FigStep [11], and VAJM [29], on MiniGPT-4. Figure 4 presents a chart comparing ASR across 13 safety-critical categories. Latent Jailbreak consistently outperforms baseline methods across all categories, demonstrating a more balanced and robust attack performance. For instance, in the critical category of Physical Harm, JailBound achieves a 97.22% ASR, significantly surpassing UMK (66.67%), which suffers from detectable pattern repetition due to uniform visual noise. FigStep, relying on converting toxic text to readable images, exhibits the lowest performance in Illegal Activity (52.58%), indicating limitations in semantic understanding. VAJM, while strong in Financial Advice (85.63%), shows inconsistency in Political Lobbying (62.75%) due to poor temporal reasoning. These results underscore that explicitly modeling the decision hyperplane enables more efficient and robust adversarial optimization.
@@ 4.4 Black-box Transferability
Figure 3: Attack Success Rate Comparison Between Our Method and Existing Methods on MM-SafetyBench Using MiniGPT-4.
@@ A.1 Extending ELK Findings to VLMs: Validating the Know-Say Disparity in VLM Settings
Inspired by research on Eliciting Latent Knowledge (ELK) on LLMs [3, 23] , recent work has revealed a fundamental insight: There exists a gap between what language models "know" internally and what they "say" in their outputs.
@@
To see whether the findings in ELK can be generized to multimodal models, we conduct the experiment on Qwen2.5-VL-7B [1], we visualize the predicted probabilities $\tilde{p}_{i}$ for both true and false samples. The results are summarized in Figure 6.

### v3

@@ Abstract
Vision-Language Models (VLMs) exhibit impressive performance, yet the integration of powerful vision encoders has significantly broadened their attack surface, rendering them increasingly susceptible to jailbreak attacks. However, lacking well-defined attack objectives, existing jailbreak methods often struggle with gradient-based strategies prone to local optima and lacking precise directional guidance, and typically decouple visual and textual modalities, thereby limiting their effectiveness by neglecting crucial cross-modal interactions. Inspired by the Eliciting Latent Knowledge (ELK) framework, we posit that VLMs encode safety-relevant information within their internal fusion-layer representations, revealing an implicit safety decision boundary in the latent space. This motivates exploiting boundary to steer model behavior. Accordingly, we propose JailBound, a novel latent space jailbreak framework comprising two stages: (1) Safety Boundary Probing, which addresses the guidance issue by approximating decision boundary within fusion layer’s latent space, thereby identifying optimal perturbation directions towards the target region; and (2) Safety Boundary Crossing, which overcomes the limitations of decoupled approaches by jointly optimizing adversarial perturbations across both image and text inputs. This latter stage employs an innovative mechanism to steer the model’s internal state towards policy-violating outputs while maintaining cross-modal semantic consistency. Extensive experiments on six diverse VLMs demonstrate JailBound’s efficacy, achieves 94.32% white-box and 67.28% black-box attack success averagely, which are 6.17% and 21.13% higher than SOTA methods, respectively. Our findings expose a overlooked safety risk in VLMs and highlight the urgent need for more robust defenses.
Warning: This paper contains potentially sensitive, harmful and offensive content.
[…]
@@ 1 Introduction
Vision Language Models (VLMs), exemplified by GPT-4o Hurst et al. (2024), LLaVA Liu et al. (2024b), and Qwen-VL Bai et al. (2025), have emerged as powerful agents capable of sophisticated multimodal reasoning. VLMs achieve their capabilities by integrating pre-trained vision encoders with Large Language Model (LLM) backbones Liu et al. (2024a), inheriting the strengths of both visual perception and natural language understanding.
Despite extensive safety alignment efforts in LLMs Lee et al. (2023); Qi et al. (2023); Ge et al. (2023); Ouyang et al. (2022) to mitigate harmful outputs, the application of similar safeguards to VLMs Liu et al. (2024d); Ding et al. (2025) remains fraught with critical challenges. The integration of visual modalities significantly broadens the attack surface, making VLMs increasingly susceptible to jailbreak attacks that threaten their safe deployment Qi et al. (2024); Gong et al. (2023); Li et al. (2025); Shayegani et al. (2023); Wang et al. (2024). Such attacks leverage adversarial perturbations, subtly introduced into either images Madry (2017); Bailey et al. (2023); Cui et al. (2024); Zhao et al. (2024) or texts Xu et al. (2024); Zou et al. (2023); Liu et al. (2023); Zhao et al. (2024); Niu et al. (2024), to elicit harmful responses. Existing jailbreak strategies, explored across diverse white-box Wang et al. (2024); Luo et al. (2024); Bailey et al. (2023); Xu et al. (2024); Cui et al. (2024) and black-box Yang et al. (2020); Chen et al. (2023, 2024); Chen and Liu (2024); Zhang et al. (2024) scenarios, highlight this vulnerability. However, these methods prove inadequate due to two primary deficiencies: 1) Depending on gradient-based optimization that are susceptible to be trapped in local optima, lacking precise directional guidance and resulting in suboptimal or overly conspicuous perturbations. 2) Processing visual and textual inputs separately, neglecting crucial cross-modal interactions.
@@
Intriguingly, recent works about the Eliciting Latent Knowledge (ELK) Burns et al. (2022); Mallen et al. (2023) on LLMs, reveal a critical phenomenon: models often encode an implicit understanding of safe versus unsafe behaviors within their internal hidden states. This suggests the existence of an internal latent safety decision boundary. Inspired by these findings, we investigate whether this concept extends to the multimodal realm of VLMs. By identifying and manipulating the latent representations associated with an internal safety boundary, we can effectively trick the model into perceiving a harmful input as acceptable for output. Building on this insight, we propose a novel attack framework, JailBound, that targets VLMs’ internal decision boundaries. JailBound first probes to approximate this latent safety boundary, and then strategically manipulates cross-modal latent representations towards the target decision, thereby eliciting policy-violating responses, as shown in Figure 1.
@@ 2 Related Work
Eliciting Latent Knowledge.  Recent research increasingly focuses on extracting and understanding the internal knowledge representations within large language models (LLMs), often called Eliciting Latent Knowledge (ELK). This inquiry is motivated by the observation that LLMs may hold knowledge suppressed by safety alignment or training biases. A foundational work is raised by Christiano et al. Paul Christiano et al. (2021), which formally posed the ELK problem. Building upon this, Burns et al. Burns et al. (2023) introduce Contrast-Consistent Search (CCS), an unsupervised method for identifying truthful directions in activation space. CCS leverages the logical consistency properties of truth to identify these directions without using any labeled data or model outputs. Mallen et al. Mallen et al. (2024) extend the ELK framework with “quirky” LLMs and datasets to evaluate the robustness of ELK methods. Our work builds upon these foundations by leveraging the insights from ELK research to address the specific challenge of jailbreaking vision-language models (VLMs). We build on ELK insights to tackle jailbreaking in VLMs, hypothesizing that policy-violating knowledge may persist internally. By adapting and extending ELK methods, we aim to exploit this latent knowledge to circumvent safety mechanisms and elicit harmful or policy-violating responses.
Jailbreak Attacks on VLMs.  The integration of vision and language modalities in VLMs introduces new attack surfaces. Recent research Qi et al. (2024); Chen et al. (2024); Gong et al. (2023); Li et al. (2025); Shayegani et al. (2023); Wang et al. (2024); Niu et al. (2024); Ma et al. (2025) have focused on jailbreak attacks, which aim to circumvent safety mechanisms and elicit harmful or policy-violating outputs from these models. For instance, FigStep Gong et al. (2023) uses typographic images to evade text-based filters, while VAJM Qi et al. (2024) shows adversarial images can bypass the safety mechanisms of VLMs, forcing universal harmful outputs. SCAV Xu et al. (2024) manipulates latent embeddings within LLMs to elicit unsafe outputs, though conceptually similar in targeting internal representations, it is restricted to uni-modal settings and lacks the ability to manipulate visual content or transfer to black-box models. Other works Bailey et al. (2023); Cui et al. (2024); Madry (2017); Zhao et al. (2024) explore adversarial perturbations, subtly modifying input images or text to mislead the VLM. These perturbations can be optimized using gradient-based methods in white-box settings Wang et al. (2024); Luo et al. (2024); Bailey et al. (2023); Xu et al. (2024); Cui et al. (2024) or through black-box approaches Yang et al. (2020); Chen et al. (2023, 2024); Chen and Liu (2024); Gong et al. (2023); Zhang et al. (2024) that rely on query feedback. Our work builds upon these advancements but is uniquely motivated by the principles of ELK, we aim to exploit this latent knowledge to craft jailbreak attacks that are both effective and semantically coherent. This approach allows us to probe the decision boundaries of VLMs, as informed by the latent representations, and generate perturbations that are more likely to succeed.
@@ 3.1 Problem Formulation
For the textual modality, An input text prompt is first processed by a tokenizer,
which converts the raw character sequence into a sequence of token IDs, $X_{t}^{\text{raw}}=[t_{1},\ldots,t_{L}]$,
where each $t_{i}\in\{1,\ldots,W\}$ and $W$ is the vocabulary size.
These token IDs are subsequently mapped to dense vector representations by then embedding layer $f_{t}:\{1,\ldots,W\}^{L}\rightarrow\mathbb{R}^{L\times D_{e}}$,
which produces the textual feature $x_{t}=f_{t}(X_{t}^{\text{raw}})\in\mathbb{R}^{L\times D_{e}}$.
For textual attacks, we append a crafted suffix $X_{t}^{\text{suffix}}$ to form $X_{t}^{\text{adv}}=[X_{t}^{\text{raw}},X_{t}^{\text{suffix}}]$, and obtain perturbed features $\tilde{x}_{t}=f_{t}(X_{t}^{\text{adv}})$.
@@ 3.4 Joint Optimization Strategy
1:Input: Dataset $M$-Safetybench $\mathcal{D}=\{(X_{v}^{(i)},X_{t}^{(i)},y^{(i)})\}_{i=1}^{N}$; VLM $F_{\theta}=(f_{v},f_{t},\phi)$; Target threshold $P_{0}$
@@
12:   Learn $(w^{(l)},b^{(l)})$ by minimizing the logistic regression loss.
@@
[…]
@@ 4.1 Experimental Setup
To validate the efficacy of our JailBound framework, we conducted extensive experiments across six Vision-Language Models (VLMs), encompassing both white-box and black-box scenarios. We leveraged the MM-SafetyBench dataset Liu et al. (2024c), a meticulously curated multimodal safety evaluation benchmark. This dataset holistically covers 13 prohibited content categories, derived from prevailing AI safety policies and emergent multimodal threat vectors. Encompassing 1,719 adversarial examples across diverse risk scenarios, each sample pairs unsafe visual content with corresponding malicious prompts. This comprehensive curation ensures a robust evaluation of our method’s resilience against both conventional and emerging threats.
@@ 4.2 Comparison With Unimodal Attacks
We first evaluate the attack efficacy of JailBound in a white-box setting across Llama-3.2-11B, Qwen2.5-VL-7B Bai et al. (2025), and MiniGPT-4 Zhu et al. (2023) in Table 2. As shown in Table 3. We use Attack Success Rate (ASR) as our evaluation metric, which refers to the proportion of successful adversarial attacks among all attempted attacks. For a more detailed discussion, please refer to Appendix. Our method demonstrates superior ASR across all models, achieving average ASRs of 94.38%, 91.40%, and 97.19%, respectively.
@@ 4.3 Comparison With Multimodal Attacks
To validate the effectiveness of our Decision Boundary-Aware Optimization, we compare JailBound against several existing attack methods, including UMKWang et al. (2024), FigStep Gong et al. (2023), and VAJM Qi et al. (2024), on MiniGPT-4. Figure 4 presents a chart comparing ASR across 13 safety-critical categories. Latent Jailbreak consistently outperforms baseline methods across all categories, demonstrating a more balanced and robust attack performance. For instance, in the critical category of Physical Harm, JailBound achieves a 97.22% ASR, significantly surpassing UMK (66.67%), which suffers from detectable pattern repetition due to uniform visual noise. FigStep, relying on converting toxic text to readable images, exhibits the lowest performance in Illegal Activity (52.58%), indicating limitations in semantic understanding. VAJM, while strong in Financial Advice (85.63%), shows inconsistency in Political Lobbying (62.75%) due to poor temporal reasoning. These results underscore that explicitly modeling the decision hyperplane enables more efficient and robust adversarial optimization.
@@ 4.4 Black-box Transferability
Figure 3: Attack Success Rate Comparison Between Our Method and Existing Methods on $M$-SafetyBench Using MiniGPT-4.
@@ A.1 Extending ELK Findings to VLMs: Validating the Know-Say Disparity in VLM Settings
Inspired by research on Eliciting Latent Knowledge (ELK) on LLMs Burns et al. (2022); Mallen et al. (2023) , recent work has revealed a fundamental insight: There exists a gap between what language models "know" internally and what they "say" in their outputs.
@@
To see whether the findings in ELK can be generized to multimodal models, we conduct the experiment on Qwen2.5-VL-7B Bai et al. (2025), we visualize the predicted probabilities $\tilde{p}_{i}$ for both true and false samples. The results are summarized in Figure 6.